OpenStudy (anonymous):

solve the system equation: 2x-4y=-4 x+2y=8 PLEASE:(

OpenStudy (anonymous):

multiply the bottom by -2, add the top and bottom together solve for y

jimthompson5910 (jim_thompson5910):

multiply both sides of equation # 2 by 2 to get 2x+4y=16 Now add this equation to the first one to get 4x=12 and solve for x to get x = 3 Now plug this into either equation to find y x+2y=8 3+2y=8 2y=5 y=5/2

OpenStudy (anonymous):

THANKS!